# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ Appendix B Main Proofs
From (B.1), we know that for any $f\in\mathcal{L}^{2}(\rho)$,
@@
B.1 Proof of the excess risk decomposition in (17)
@@
B.2 Proof of Theorem 2: bounding the excess risk
@@
$\mathbb{E}[\ell_{f}(Y,X)]$ $=\mathbb{E}[h_{f}(Z,X)]$ by (B)
$\leq 2\mathbb{E}_{n}[h_{f}(Z,X)]+C\Lambda+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ by ($\text{by }\mathcal{E}(h)$
@@
$\qquad+C\Lambda+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ by (B)
@@
$\mathbb{E}_{n}[\ell_{f_{\widehat{\alpha}}}(Y,X)]+\lambda\|f_{\widehat{\alpha}}\|^{2}_{K}-\lambda\|f_{\mathcal{H}}\|_{K}^{2}~{}\leq~{}\widehat{\alpha}\left\{\mathbb{E}_{n}[\ell_{\widehat{f}}(Y,X)]+\lambda\|\widehat{f}\|^{2}_{K}-\lambda\|f_{\mathcal{H}}\|_{K}^{2}\right\}.$
@@
then (77) also holds for $\widehat{f}$ in lieu of $f_{\widehat{\alpha}}$, which contradicts (71). To prove (77), recalling that $f_{\widehat{\alpha}}\in\uline{\mathcal{F}_{b}}$, invoking the uniform bound in (B.2) with $f=f_{\widehat{\alpha}}$ gives
@@
$-\mathbb{E}[\ell_{f_{\widetilde{\alpha}}}(Y,X)]$ $=-\mathbb{E}[h_{f_{\widetilde{\alpha}}}(Z,X)]$ by (B)
@@
Meanwhile, since $f_{\widetilde{\alpha}}\in\uline{\mathcal{F}_{b}}$, invoking (B.2) with $f=f_{\widetilde{\alpha}}$ gives
@@
$\leq 2\|f_{\mathcal{H}}\|_{K}^{2}~{}\mathbb{E}\Delta_{\widehat{g}}+2\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}+4\lambda\|f_{\mathcal{H}}\|^{2}_{K}-2\lambda\|f_{\widetilde{\alpha}}\|^{2}_{K}$ by (B.2)
@@
[…]
@@
B.2.1 Uniform upper bounds of $\mathbb{E}[h_{f}(Z,X)]$ over $f\in\mathcal{F}_{b}$
@@
^${\widehat{\mathcal{R}}_{n}}(\mathcal{F}):=\mathbb{E}_{\varepsilon}\Big{[}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f\left({v}_{i}\right)\Big{]},$ (82)
@@
The sub-root function is an essential ingredient for establishing fast rates in statistical learning problems. Precisely, when deriving the bound on the error of a learning algorithm, the key step is to select an appropriate sub-root function, and its corresponding fixed point will appear in the prediction error bound. In Appendix C.1 we review the definition and some useful properties of sub-root functions.
@@
This shows the function class $\mathcal{F}_{h}$ is uniformly bounded within $[-V,V]$.
@@
where the last step uses (3.1) with $\theta=1$, we have
@@
$\leq V\left(\mathbb{E}[h_{f}(Z,X)]+4\|f_{\mathcal{H}}\|_{K}^{2}\mathbb{E}\Delta_{\widehat{g}}+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}\right)$ by (3.1) with $\theta=1$
=(87)${=}T(h_{f}).$ ) .
@@
Note that (B.2.1) ensures that $\psi_{j}(\theta_{j})$ is $(8\kappa\|f_{\mathcal{H}}\|_{K}+2\|f^{*}\|_{\infty})$-Lipschitiz. Invoking Lemma 23 yields
@@
so that, together with (B.2.1),
@@
To find a sub-root function that meets the requirement (91), recall that the function $\psi_{x}(\cdot)$ in (86) is sub-root. It follows from (B.2.1) that
@@
Finally, the proof is completed by invoking Lemma 22 with $-a=b=B=V$, $L=4\|f_{\mathcal{H}}\|_{K}^{2}\mathbb{E}\Delta_{\widehat{g}}+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ and $\theta=2$. ∎
B.2.2 Bounding from above the fixed point of the local Rademacher complexity $\psi_{x}(\delta)$ by that of the kernel complexity $R_{x}(\delta)$
This section is devoted to deriving an upper bound of $\delta_{x}$ which is the fixed point of $\psi_{x}(\cdot)$ in (86) via a series of useful lemmas in Section B.3, followed by a corollary of Lemma 3. Recall that $\Delta_{\widehat{g}}$ from (18). Further recall $\delta_{n,\widehat{g}}$ and $V_{K}$ from (67) and (68), respectively.
@@
Recalling ^${\widehat{\mathcal{R}}_{n}}$ from (82), we first link $\psi_{x}(\delta)$ to its following data-dependent version
@@
B.2.3 Local uniform upper bounds of the cross-term
@@
:=$\mathcal{E}:=\left\{\beta\in\mathbb{R}^{\infty}:\sum_{j=1}^{\infty}\beta_{j}^{2}{\nu}_{j}\leq 1+9\|f_{\mathcal{H}}\|_{K}^{2}\right\},$ (112)
@@
$[H]_{ab}=\sum_{j=1}^{\infty}\frac{1}{\nu_{j}}(\widehat{\phi}_{x,j}\circ\widehat{g})(X_{a})(\widehat{\phi}_{x,j}\circ\widehat{g})(X_{b}),\qquad\forall~{}a,b\in[n].$
@@
Below we apply Lemma 22 to the function class $\Xi^{2}_{b}$ with $T=T,L=0$ and $B=-a=b=V$.
Step 1: Verification of the boundedness of $\Xi^{2}_{b}$. By using (89), the function class $\Xi^{2}_{b}$ is uniformly bounded with range $[-V,V]$.
@@
which implies that $\xi_{f}^{2}(X_{i})$ is $(8\kappa\|f_{\mathcal{H}}\|_{K})$-Lipschitz in $\xi_{f}(X_{i})$ for all $i\in[n]$. Then, by repeating the similar arguments of proving (B.2.1), we can apply Ledoux–Talagrand contraction inequality in Lemma 23 to obtain
@@
By using (3.1) with $\theta=1$, we have
@@
The proof can be done by largely following the argument used in the proof of Lemma 6.6 in Bartlett et al. (2005). First, it is easy to see that the non-zero eigenvalues of $\widehat{T}_{x,K}$ are at most $n$ in number because the range of $\widehat{T}_{x,K}$ is at most $n$-dimensional.
@@
implies that $\mu$ is also an eigenvalue of $\mathbf{K}_{x}$. On the other hand, suppose that $v$ is an non-trivial eigenvector (that is, $\|\mathbf{K}_{x}v\|_{2}^{2}>0$) of $\mathbf{K}_{x}$ with non-zero eigenvalue $\mu$. Then by letting $f_{v}:=n^{-1/2}\sum_{j=1}^{n}v_{j}K(\widehat{Z}_{j},\cdot)$, we find that
@@
implying that either $f_{v}=0$ or $\mu$ is an eigenvalue of $\mathbf{K}_{x}$. The former is impossible since
@@
B.3 Technical lemmas for proving Lemma 4
@@
B.3.1 Relating local Rademacher complexity to its empirical counterpart
@@
Let $\mathcal{F}$ be a function class with ranges in $[-G,G]$ and $U_{1},\ldots,U_{n}$ be the i.i.d. copies of some random variable $U$. Fix any $t\geq 0$. For any $\delta$ such that
@@
B.3.2 Relating empirical local Rademacher complexity in (41) to empirical kernel complexity $\widehat{R}_{x}(\delta)$
@@
^${\widehat{\mathcal{R}}_{n}}\left(\big{\{}\xi_{f}\in\Xi_{b}:~{}\mathbb{E}_{n}[\xi^{2}_{f}(X)]\leq\delta\big{\}}\right)$ (126)
@@
^${\widehat{\mathcal{R}}_{n}}\left(\big{\{}\xi_{f}\in\Xi_{b}:~{}\mathbb{E}_{n}[\xi^{2}_{f}(X)]\leq\delta\big{\}}\right)\leq\sqrt{10}~{}(\|f_{\mathcal{H}}\|_{K}\vee 1)~{}\widehat{R}_{x}(\delta).$
@@
B.3.3 Relating $\widehat{R}_{x}(\delta)$ to $\widehat{R}(\delta)$
@@
$\begin{split}&\Sigma_{x}:\ell^{2}(\mathbb{N})\to\ell^{2}(\mathbb{N}),\qquad\Sigma_{x}=\frac{1}{n}\sum_{i=1}^{n}\Phi(\widehat{Z}_{i})\Phi(\widehat{Z}_{i})^{\top};\\ &\Sigma:\ell^{2}(\mathbb{N})\to\ell^{2}(\mathbb{N}),\qquad~{}~{}\Sigma=\frac{1}{n}\sum_{i=1}^{n}\Phi(Z_{i})\Phi(Z_{i})^{\top}.\\ \end{split}$ (135)
@@
$\begin{split}&A(\delta)=(\Sigma+\delta I)^{-\frac{1}{2}}(\Sigma-\Sigma_{x})(\Sigma+\delta I)^{-\frac{1}{2}};\\ &B(\delta)=(\Sigma+\delta I)^{-1}(\Sigma-\Sigma_{x})(\Sigma+\delta I)^{-1}.\\ \end{split}$
@@
[…]
@@
B.3.4 Relating $\widehat{R}(\delta)$ to $R(\delta)$

### v2

@@ Appendix A Main Proofs
From (A.1), we know that for any $f\in\mathcal{L}^{2}(\rho)$,
@@
A.1 Proof of (17) and the excess risk decomposition in (16)
@@
Proof.
We have that for any $\theta\geq 1$, the elementary inequality $(a+b)^{2}\leq xa^{2}+b^{2}/x$ for $x>0$, so that
$~=~(1+\theta)~\mathbb{E}\left\langle f_{\mathcal{H}},K_{Z}-K_{\widehat{g}(X)}\right\rangle_{K}^{2}+{1+\theta\over\theta}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$
$~\leq~(1+\theta)\|f_{\mathcal{H}}\|_{K}^{2}~\mathbb{E}\Delta_{\widehat{g}}+{1+\theta\over\theta}\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}.$
The second step uses the reproducing property. ∎
A.2 Proof of Theorem 1: bounding the excess risk
@@
$\mathbb{E}[\ell_{f}(Y,X)]$ $=\mathbb{E}[h_{f}(Z,X)]$ by (A)
$\leq 2\mathbb{E}_{n}[h_{f}(Z,X)]+C\Lambda+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ $\text{by }\mathcal{E}(h)$
@@
$\qquad+C\Lambda+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ by (A)
@@
[…]
@@
then (70) also holds for $\widehat{f}$ in lieu of $f_{\widehat{\alpha}}$, which contradicts (64). To prove (70), recalling that $f_{\widehat{\alpha}}\in\uline{\mathcal{F}_{b}}$, invoking the uniform bound in (A.2) with $f=f_{\widehat{\alpha}}$ gives
@@
$-\mathbb{E}[\ell_{f_{\widetilde{\alpha}}}(Y,X)]$ $=-\mathbb{E}[h_{f_{\widetilde{\alpha}}}(Z,X)]$ by (A)
@@
Meanwhile, since $f_{\widetilde{\alpha}}\in\uline{\mathcal{F}_{b}}$, invoking (A.2) with $f=f_{\widetilde{\alpha}}$ gives
@@
$\leq 2\|f_{\mathcal{H}}\|_{K}^{2}~\mathbb{E}\Delta_{\widehat{g}}+2\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}+4\lambda\|f_{\mathcal{H}}\|^{2}_{K}-2\lambda\|f_{\widetilde{\alpha}}\|^{2}_{K}$ by (A.2)
@@
$\leq\alpha\|f\|_{K}^{2}+(1-\alpha)\|f_{\mathcal{H}}\|^{2}_{K}$ $\text{by the convexity of the quadratic function}.$
@@
A.2.1 Uniform upper bounds of $\mathbb{E}[h_{f}(Z,X)]$ over $f\in\mathcal{F}_{b}$
@@
${\widehat{\mathcal{R}}_{n}}(\mathcal{F}):=\mathbb{E}_{\varepsilon}\Big[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f\left({v}_{i}\right)\Big],$ (75)
@@
The sub-root function is an essential ingredient for establishing fast rates in statistical learning problems. Precisely, when deriving the bound on the error of a learning algorithm, the key step is to select an appropriate sub-root function, and its corresponding fixed point will appear in the prediction error bound. In Appendix B.1 we review the definition and some useful properties of sub-root functions.
@@
This shows the function class $\mathcal{F}_{h}$ is uniformly bounded within [V,V][-V,V].
@@
where the last step uses (17) with $\theta=1$, we have
@@
$\leq V\left(\mathbb{E}[h_{f}(Z,X)]+4\|f_{\mathcal{H}}\|_{K}^{2}\mathbb{E}\Delta_{\widehat{g}}+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}\right)$ by (17) with $\theta=1$
[…]
@@
Note that (A.2.1) ensures that $\psi_{j}(\theta_{j})$ is $(8\kappa\|f_{\mathcal{H}}\|_{K}+2\|f^{*}\|_{\infty})$-Lipschitiz. Invoking Lemma 23 yields
@@
so that, together with (A.2.1),
@@
To find a sub-root function that meets the requirement (84), recall that the function $\psi_{x}(\cdot)$ in (79) is sub-root. It follows from (A.2.1) that
@@
Finally, the proof is completed by invoking Lemma 22 with a=b=B=V-a=b=B=V, $L=4\|f_{\mathcal{H}}\|_{K}^{2}\mathbb{E}\Delta_{\widehat{g}}+4\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ and $\theta=2$. ∎
A.2.2 Bounding from above the fixed point of the local Rademacher complexity $\psi_{x}(\delta)$ by that of the kernel complexity $R(\delta)$
This section is devoted to deriving an upper bound of $\delta_{x}$ which is the fixed point of $\psi_{x}(\cdot)$ in (79) via a series of useful lemmas in Section A.3, followed by a corollary of Lemma 3. Recall that $\Delta_{\widehat{g}}$ from (18). Further recall $\delta_{n,\widehat{g}}$ and $V_{K}$ from (60) and (61), respectively.
@@
Recalling $\widehat{\mathcal{R}}_{n}$} from (75), we first link $\psi_{x}(\delta)$ to its following data-dependent version
@@
A.2.3 Local uniform upper bounds of the cross-term
@@
$\mathcal{E}:=\left\{\beta\in\mathbb{R}^{\infty}:\sum_{j=1}^{\infty}\beta_{j}^{2}{\nu}_{j}\leq 1+9\|f_{\mathcal{H}}\|_{K}^{2}\right\},$ (105)
@@
$H]_{ab}=\sum_{j=1}^{\infty}\frac{1}{\nu_{j}}(\widehat{\phi}_{x,j}\circ\widehat{g})(X_{a})(\widehat{\phi}_{x,j}\circ\widehat{g})(X_{b}),\qquad\forall~a,b\in[n].$
@@
Below we apply Lemma 22 to the function class $\Xi^{2}_{b}$ with $T=T,L=0$ and B=a=b=VB=-a=b=V.
Step 1: Verification of the boundedness of $\Xi^{2}_{b}$. By using (82), the function class $\Xi^{2}_{b}$ is uniformly bounded with range [V,V][-V,V].
@@
which implies that $\xi_{f}^{2}(X_{i})$ is $(8\kappa\|f_{\mathcal{H}}\|_{K})$-Lipschitz in $\xi_{f}(X_{i})$ for all $i\in[n]$. Then, by repeating the similar arguments of proving (A.2.1), we can apply Ledoux–Talagrand contraction inequality in Lemma 23 to obtain
@@
By using (17) with $\theta=1$, we have
@@
The proof can be done by largely following the argument used in the proof of Lemma 6.6. in Bartlett et al. (2005). First, it is easy to see that the non-zero eigenvalues of $\widehat{T}_{x,K}$ are at most $n$ in number because the range of $\widehat{T}_{x,K}$ is at most $n$-dimensional.
@@
implying that $\mathbf{K}_{x}f_{n}=\mu f_{n}.$ Together with
[…]
implies that $\mu$ is also an eigenvalue of $\mathbf{K}_{x}$. On the other hand, suppose that $v$ is an non-trivial eigenvector (that is, $v^{\top}\mathbf{K}_{x}v>0$) of $\mathbf{K}_{x}$ with non-zero eigenvalue $\mu$. Then by letting $f_{v}:=n^{-1/2}\sum_{j=1}^{n}v_{j}K(\widehat{Z}_{j},\cdot)$, we find that
@@
implying that either $\langle f_{v},f_{v}\rangle_{K}=0$ or $\mu$ is an eigenvalue of $\widehat{T}_{x,K}$. The former is impossible since
@@
A.3 Technical lemmas for proving Lemma 4
@@
A.3.1 Relating local Rademacher complexity to its empirical counterpart
@@
Let $\mathcal{F}$ be a function class with ranges in [G,G][-G,G] and $U_{1},\ldots,U_{n}$ be the i.i.d. copies of some random variable $U$. Fix any $t\geq 0$. For any $\delta$ such that
@@
A.3.2 Relating empirical local Rademacher complexity in (36) to empirical kernel complexity $\widehat{R}_{x}(\delta)$
@@
${\widehat{\mathcal{R}}_{n}}\left(\big\{\xi_{f}\in\Xi_{b}:~\mathbb{E}_{n}[\xi^{2}_{f}(X)]\leq\delta\big\}\right)$ (119)
@@
${\widehat{\mathcal{R}}_{n}}\left(\big\{\xi_{f}\in\Xi_{b}:~\mathbb{E}_{n}[\xi^{2}_{f}(X)]\leq\delta\big\}\right)\leq\sqrt{10}~(\|f_{\mathcal{H}}\|_{K}\vee 1)~\widehat{R}_{x}(\delta).$
@@
A.3.3 Relating $\widehat{R}_{x}(\delta)$ to $\widehat{R}(\delta)$
@@
[…]
@@
[…]
@@
$\leq~$ $\frac{3}{2}\mathbb{E}\Delta_{\widehat{g}}+\frac{28\kappa^{2}\log(2/\eta)}{3n}.$
@@
A.3.4 Relating $\widehat{R}(\delta)$ to $R(\delta)$